# A NEW POLYNOMIAL FOR CHECKERBOARD-COLORABLE 4-VALENT VIRTUAL GRAPHS

### v2

@@ 2.3 The Jones-Kauffman polynomial of a virtual link
The Kauffman bracket polynomial is a function from the set of unoriented virtual link diagrams in the oriented plane to the ring of Laurent polynomials with integer coefficients in an indeterminate $q$. It maps a virtual link $L$ to $\left\langle L\right\rangle\in\mathbb{Z}[q^{-1},q]$ and it is defined by the following relations:
@@ 3 The main results and an application
We say that a $2$-diagraph is vertex-signed if each vertex is equipped with a sign $+$ or --.
@@
Let $G$ be a $2$-diagraph with $n$ vertices $v_{i}$, $1\leq i\leq n$, and $\gamma$ be an Euler circuit of $G$. At each vertex $v_{i}$, we denote the two incoming edges by $e_{i_{1}}$ and $e_{i_{2}}$, and we denote the two outgoing edges by $e_{i_{3}}$ and $e_{i_{4}}$. We assume that at a fixed vertex $v_{i}$ the Euler circuit $\gamma$ of $G$ traverses successively the edges $e_{i_{1}}$ and $e_{i_{3}}$ in the first visit, and then it traverses successively $e_{i_{2}}$ and $e_{i_{4}}$ in the second visit. Merge the edges $e_{i_{1}}$ and $e_{i_{3}}$, i.e., replace them with a single new edge going from the tail of $e_{i_{1}}$ to the head of $e_{i_{3}}$. Also, we merge the edges $e_{i_{2}}$ and $e_{i_{4}}$ using the same technique and we delete the vertex $v_{i}$ to get a new graph with one less vertex. In the vicinity of the deleted vertex $v_{i}$, put a marker connecting the traces of the deleted vertex on each one of the new edges. We label each marker by $A$, $a$, $B$, or $b$ according to the coloring, the sign of the vertex and the order of the edges when traversing the path as explained in Figure 9. Such labeled marker for the vertex $v_{i}$ is denoted by “$i$-th marker”. We iterate this merging process until all the vertices are deleted and we obtain a single immersed circle in the plane with $n$ labeled markers in the vicinities of the deleted vertices. We call such data the $\gamma$-state of the corresponding Euler circuit $\gamma$ of $G$. In Figure 10, we depict the $\gamma$-state of the Euler circuit $\gamma$ given in Figure 4.
@@
We enclose this subsection by stating the following theorem whose proof will be given in the next section which implies that the polynomial $X_{G}(q)$ satisfies a skein relation. For this end, we let $G^{v}$ be a checkerboard-colorable vertex-signed $2$-digraph with a fixed vertex $v$ depicted in Figure 11. The digraphs resulting from the two possible merging operations at $v$ described in Section 3 will be denoted by $G_{0}^{v}$ and $G_{1}^{v}$ respectively according to the scheme in Figure 11.
@@
Let $L$ be an oriented checkerboard-colorable virtual link. Let $G$ be the shadow graph of a checkerboard-colorable diagram of $L$, then $f_{L}(q)=(-q)^{-3\omega(L)}X_{G}(q)$, where $f_{L}(q)$ and $\omega(L)$ are the Jones-Kauffman polynomial and the writhe of the oriented virtual link $L$, respectively.
@@ 4 Proof of the Main Theorems
As an example, in Figure 4 $\gamma$ can be written as $v_{1}e_{1}v_{2}e_{2}v_{4}e_{3}v_{1}e_{4}v_{2}e_{5}v_{3}e_{6}v_{4}e_{7}v_{3}e_{8}v_{1}$ while $\gamma^{v_{3}v_{4}}$ can be written as $v_{1}e_{1}v_{2}e_{2}v_{4}e_{3}v_{3}e_{4}v_{4}e_{5}v_{1}e_{6}v_{2}e_{7}v_{3}e_{8}v_{1}$. The $\gamma$-state is depicted in Figure 10 while the $\gamma^{v_{3}v_{4}}$-state is depicted in Figure 14. One can observe that the $3$rd marker and the $4$th marker are the only ones which change labels when one compares the two states.
@@
According to [1, Lemma 4], any circuit of $G$ can be obtained from the circuit $\gamma$ by transposing on some interlaced pair of vertices.
@@
    $\Leftrightarrow$ there is an $\ \text{there is an edge}\ e\ \text{connecting}\ v_{k}\ \text{to}\ v_{j}\ \text{in}\ H(\gamma^{v_{i}v_{j}})\ \text{and}\ k\neq i$
    $\Leftrightarrow$ there is an $\ \text{there is an edge}\ e\ \text{connecting}\ v_{k}\ \text{to}\ v_{i}\ \text{in}\ H(\gamma^{v_{i}v_{j}})_{v_{i}v_{j}}\ \text{and}\ k\neq j$
    $\Leftrightarrow$ there is an edge e $\ \text{there is an edge $e$ in}\ H^{v_{i}v_{j}}\ \text{connecting}\ v_{k}\ \text{to}\ v_{i}\ \text{and}\ k\neq j$
    $\Leftrightarrow$ there is an edge e $\ \text{there is an edge $e$ in}\ H\ \text{connecting}\ v_{k}\ \text{to}\ v_{i}\ \text{and}\ k\neq j$
@@
    $\Leftrightarrow$ there is an $\ \text{there is an edge}\ e\ \text{connecting}\ v_{k}\ \text{to}\ v_{j}\ \text{in}\ H(\gamma^{v_{i}v_{j}})=(H^{v_{i}v_{j}}(\gamma))_{v_{i}v_{j}}$
    $\Leftrightarrow$ there is an $\ \text{there is an edge}\ e\ \text{connecting}\ v_{k}\ \text{to}\ v_{i}\ \text{in}\ H^{v_{i}v_{j}}$
@@
    $\Leftrightarrow$ there is an $\ \text{there is an edge}\ e\ \text{connecting}\ v_{h}\ \text{to}\ v_{k}\ \text{in}\ H(\gamma^{v_{i}v_{j}})_{v_{i}v_{j}}$
    $\Leftrightarrow$ there is an $\ \text{there is an edge}\ e\ \text{connects}\ v_{h}\ \text{to}\ v_{k}\ \text{in}$
@@
    $\Leftrightarrow$ there is an $\ \text{there is an edge}\ e\ \text{connecting}\ v_{h}\ \text{to}\ v_{k}\ \text{in}\ H$
@@
      In Table 4, we state the $i$-th and $(i+1)$-th letters of the activity words of the four circuits in all possible cases using Remark 4.2 and Remark 4.5. In the eight cases marked by an ”$\times$” in the table, we have

### v3

@@ 2.3 The Jones-Kauffman polynomial of a virtual link
The Kauffman bracket polynomial is a function from the set of unoriented virtual link diagrams in the oriented plane to the ring of Laurent polynomials with integer coefficients in an indeterminate $q$. It maps a virtual link diagram $L$ to $\left\langle L\right\rangle\in\mathbb{Z}[q^{-1},q]$ and it is defined by the following relations:
@@ 3 The main results and an application
We say that a $2$-digraph is vertex-signed if each vertex is equipped with a sign ++ or -.
@@
Let $G$ be a $2$-digraph with $n$ vertices $v_{i}$, $1\leq i\leq n$, and $\gamma$ be an Euler circuit of $G$. At each vertex $v_{i}$, we denote the two incoming edges by $e_{i_{1}}$ and $e_{i_{2}}$, and we denote the two outgoing edges by $e_{i_{3}}$ and $e_{i_{4}}$. We assume that at a fixed vertex $v_{i}$ the Euler circuit $\gamma$ of $G$ traverses successively the edges $e_{i_{1}}$ and $e_{i_{3}}$ in the first visit, and then it traverses successively $e_{i_{2}}$ and $e_{i_{4}}$ in the second visit. Merge the edges $e_{i_{1}}$ and $e_{i_{3}}$, i.e., replace them with a single new edge going from the tail of $e_{i_{1}}$ to the head of $e_{i_{3}}$. Also, we merge the edges $e_{i_{2}}$ and $e_{i_{4}}$ using the same technique and we delete the vertex $v_{i}$ to get a new graph with one less vertex. In the vicinity of the deleted vertex $v_{i}$, put a marker connecting the traces of the deleted vertex on each one of the new edges. We label each marker by $A$, $a$, $B$, or $b$ according to the coloring, the sign of the vertex and the order of the edges when traversing the path as explained in Figure 9. Such labeled marker for the vertex $v_{i}$ is denoted by “$i$-th marker”. We iterate this merging process until all the vertices are deleted and we obtain a single immersed circle in the plane with $n$ labeled markers in the vicinities of the deleted vertices. We call such data the $\gamma$-state of the corresponding Euler circuit $\gamma$ of $G$. In Figure 10, we depict the $\gamma$-state of the Euler circuit $\gamma$ given in Figure 4.
@@
We enclose this subsection by stating the following theorem whose proof will be given in the next section which implies that the polynomial $X_{G}(q)$ satisfies a skein relation. For this end, we let $G^{v}$ be a checkerboard-colorable vertex-signed $2$-digraph with a fixed vertex $v$ depicted in Figure 11. The $2$-digraphs resulting from the two possible merging operations at $v$ described in Section 3 will be denoted by $G_{0}^{v}$ and $G_{1}^{v}$ respectively according to the scheme in Figure 11.
@@
Let $L$ be an oriented checkerboard-colorable virtual link. Let $G$ be the shadow graph of a checkerboard-colorable diagram of $L$ made into a vertex-signed $4$-valent graph, then $f_{L}(q)=(-q)^{-3\omega(L)}X_{G}(q)$, where $f_{L}(q)$ and $\omega(L)$ are the Jones-Kauffman polynomial and the writhe of the oriented virtual link $L$, respectively.
@@ 4 Proof of the Main Theorems
As an example, in Figure 4 $\gamma$ can be written as $v_{1}e_{1}v_{2}e_{2}v_{4}e_{3}v_{1}e_{4}v_{2}e_{5}v_{3}e_{6}v_{4}e_{7}v_{3}e_{8}v_{1}$ while $\gamma^{v_{3}v_{4}}$ can be written as $v_{1}e_{1}v_{2}e_{2}v_{4}e_{7}v_{3}e_{6}v_{4}e_{3}v_{1}e_{4}v_{2}e_{5}v_{3}e_{8}v_{1}$. The $\gamma$-state is depicted in Figure 10 while the $\gamma^{v_{3}v_{4}}$-state is depicted in Figure 14. One can observe that the 33rd marker and the $4$th marker are the only ones which change labels when one compares the two states.
@@
According to [1, Lemma 4], any circuit of $G$ can be obtained from the circuit $\gamma$ by transposing on some interlaced pairs of vertices.
@@
    $\Leftrightarrow$ $\ \text{there is an edge}\ e\ \text{connecting}\ v_{k}\ \text{to}\ v_{j}\ \text{in}\ H(\gamma^{v_{i}v_{j}})\ \text{and}\ k\neq i$
    $\Leftrightarrow$ $\ \text{there is an edge}\ e\ \text{connecting}\ v_{k}\ \text{to}\ v_{i}\ \text{in}\ H(\gamma^{v_{i}v_{j}})_{v_{i}v_{j}}\ \text{and}\ k\neq j$
    $\Leftrightarrow$ $\ \text{there is an edge $e$ in}\ H^{v_{i}v_{j}}\ \text{connecting}\ v_{k}\ \text{to}\ v_{i}\ \text{and}\ k\neq j$
    $\Leftrightarrow$ $\ \text{there is an edge $e$ in}\ H\ \text{connecting}\ v_{k}\ \text{to}\ v_{i}\ \text{and}\ k\neq j$
@@
    $\Leftrightarrow$ $\ \text{there is an edge}\ e\ \text{connecting}\ v_{k}\ \text{to}\ v_{j}\ \text{in}\ H(\gamma^{v_{i}v_{j}})=(H^{v_{i}v_{j}}(\gamma))_{v_{i}v_{j}}$
    $\Leftrightarrow$ $\ \text{there is an edge}\ e\ \text{connecting}\ v_{k}\ \text{to}\ v_{i}\ \text{in}\ H^{v_{i}v_{j}}$
@@
    $\Leftrightarrow$ $\ \text{there is an edge}\ e\ \text{connecting}\ v_{h}\ \text{to}\ v_{k}\ \text{in}\ H(\gamma^{v_{i}v_{j}})_{v_{i}v_{j}}$
    $\Leftrightarrow$ $\ \text{there is an edge}\ e\ \text{connects}\ v_{h}\ \text{to}\ v_{k}\ \text{in}$
@@
    $\Leftrightarrow$ $\ \text{there is an edge}\ e\ \text{connecting}\ v_{h}\ \text{to}\ v_{k}\ \text{in}\ H$
@@
      In Table 4, we state the $i$-th and $(i+1)$-th letters of the activity words of the four circuits in all possible cases using Remark 4.2 and Remark 4.5. In the eight cases marked by an $\times$” in the table, we have